OpenStudy (anonymous):

Josh has 4 shirts white, blue, green, and red and 4 pairs of socks in the same colors. Suppose Josh grabs a shirt and a pair of socks without looking. What is the probability that his shirt and socks won't match?

OpenStudy (anonymous):

P = 1 - P(same color for both shirt and socks)

OpenStudy (anonymous):

P = 1 - 4/(12C3) = 54/55